A car is stopped at the traffic light. It then travels along a straight road so that its distance from the light is given by π‘₯(𝑑) = 𝑏𝑑2 βˆ’ 𝑐𝑑3 where b=2.40 m/s2 and c=0.120 m/s3.

a) Calculate the average velocity of the car for the time interval t=0 to t=10.0 s.

b) Calculate the instantaneous velocity of the car at t=0 and t=10.0s

c) How long after starting again from rest is the car again at rest.


1
Expert's answer
2022-03-02T10:07:18-0500

a)

"v_a=\frac x{\Delta t}=-20~\frac ms,"

b)

"v_i=2bt-3ct^2,"

"v_i(0)=0~\frac ms,"

"v_i(10)=12~\frac ms,"

c)

"t=\frac bc=20~s."